June 25, 1999 (Friday)

SESSION Computer Algebra meets Education I Bernhard Kutzler

A Approximate Algebraic Computation: towards Symbolic-Numeric Algorithms Tateaki Sasaki Matu-Tarow Noda Bernard Mourrain & Robert Corless

B Combinatorial and Computational Methods in Algebraic Geometry Jie-Tai Yu & Vladimir Shpilrain

C Computer Algebra for Dynamical Systems and Mechanics I Victor Edneral Ilias Kotsireas & Nikolay Vasiliev

D Applications of Computer Algebra to Signal Processing Jeremy Johnson & Markus Pueschel

E

9:00 am Setting the Tone: CAS as pedagogical tools Bernhard Kutzler New results on the Table Maker’s Dilemma Jean-Michel Muller & Vincent Lefevre* Explicit computations for polynomial maps Pierrette Casson-Nogus Computer Algebraic Approximations— Creation of Approximated Solutions of Scientific and Engineering Problems in Symbolic Form Victor Edneral Minimal Syzygies and Multidimensional Filter Design Hyungju Park

9:30 am Basic Skills versus Technology—not a Contradiction but a Completion Josef Boehm Polynomial Algebra with Coefficients of Limited Accuracy Hans Stetter The Parameterization Problem for Algebraic Surfaces Josef Schicho Group Representations and Automatic Derivation of Fast Signal Transforms Markus Pueschel

10:00 am The Use of Mathematics Specific Technology in Capstone Mathematics Teacher Preparation Courses Gary A. Harris A System of Automatic Algorithm Stabilization Hiroshi Sekigawa Computing subfields in pure trascendental extensions Jaime Gutierrez & Rosario Rubio Solving the problem of stabilization of a gyroscopic system with the help of Computer Algebra A. V. Banshchikov, L. A. Bourlakova & V. D. Irtegov A Wreath Product Approach to Signal and Image Processing Dan Rockmore

10:30 am Some Reflections About the Impact of Computer Algebra Systems in the Ordering of the Curricula of Mathematics Justo Cabezas & Eugenio Roanes-Lozano* To Be Announced Hiroshi Kai Again x + x^{2}y + z^{2} + t^{3} = 0 Lenny Makar-Limanov About investigation of systems with first integrals L. A. Bourlakova & V. D. Irtegov Groebner Bases and Wavelet Design Ivan Selesnick

11:00 am  BREAK

11:30 am A Computer Classroom Lecture: Asymptotic Convergence of Rational Functions Bill Pletsch To Be Announced Victor Y. Pan Invariants of algebraic varieties Vladimir Shpilrain Analytical Approach for Simplifying Dynamical Systems of Polynomial Type J. Palacin & P. Yanguas To Be Announced Jean-Charles Faugere & Fabrice Rouillier

12:00 pm Exam questions when using CAS for school mathematics teaching Vlasta Kokol-Voljc Mechanism of Cancellation Errors in Multivariate Hensel Construction with Floating-point Numbers Tateaki Sasaki Communications of the Poisson Series Processors PSPC with General Scientific Software A. Abad, A. Gavin & J. F. San-Juan To Be Announced Jeremy Johnson

12:30 pm Using the TI-92Plus: Examples Michel Beaudin Exacts predicates for arrangement of arcs of circle A. Fronville Parallelisation of Perturbation Analysis: the method of Multiple Scales applied to vibration problems Raya Khanin & Matthew Cartmell

1:00 pm Why use CAS with TI89 in mathematical education? Classrooms experiments. Bengt Ahlander Approximate p-adic GCD computation D. Rupprecht Some computational experiments in Celestial Mechanics Jean-Charles Faugere

June 25, 1999 (Friday)

SESSION Computer Algebra meets Education I Bernhard Kutzler

A Approximate Algebraic Computation: towards Symbolic-Numeric Algorithms Tateaki Sasaki Matu-Tarow Noda Bernard Mourrain & Robert Corless

B Toric Ideals and Integer Programming Lorenzo Robbiano

C Computer Algebra for Dynamical Systems and Mechanics I Victor Edneral Ilias Kotsireas & Nikolay Vasiliev

D Computer Algebra in the Spanish Educational System Lola Rodriguez

E

3:30 pm Using Logo as a Scaffolding to Develop Algebraic Thinking in a Virtual Classroom Garry Clark* & Edward Redden Towards numerical stability in normal form algorithms B. Mourrain Computing Toric Ideals Lorenzo Robbiano Some remarks on differential Hilbert polynomials in two variables G. Carra Ferro Aparicion de nuevos contenidos curriculares en matematicas mediante la aplicacion de nuevas tecnologias Justo Cabezas Corchero

4:00 pm Discovery Experiences in Synthetic Geometry with DERIVE Miguel de Guzman Hybrid Rational Function Approximation and its Applications Matu T. Noda Extended Characteristic Sets of Finitely Generated Differential Ideals G. Carra Ferro & Vladimir Gerdt Las herramientas euclideas y el ordenador Tomas Recio

4:30 pm About how to implement Geometric Transformations on a CAS E. Roanes-Macias & E. Roanes-Lozano* Approximate polynomial decomposition M. Giesbrecht New algorithms for Computing Toric Ideals Anna Bigatti Enumeration of intersections of Newton’s polyhedrons Alexander Aranson Matematicas en Pantalla Rafael Perez Gomez & Miguel Posadas

5:00 pm  BREAK

5:30 pm A number-based approach to introductory algebra: A study with 11-12 year olds using graphic calculators Tenoch E. Cedillo A.

On approximate ideals B. Trager or P. Gianni Decomposing Graver Test Sets in Stochastic Programming Raymond Hemmecke Symbolic Computation of Formal Solutions for 2 and 3 Dimensional Dynamical Systems G. Eichenmueller Diferentes opciones para la resolucion de problemas con calculadoras graficas Agustin Carrillo

6:00 pm Instrumentation processes underlying the use of TI92 at high school level Badr Defouad

Initial  Ideals  of

Toric Ideals and Group Relaxations in Integer Programming Serkan Hosten A new classification of planar homogeneous quadratic systems Driss Boularas Cambios curriculares en la enseñanza del Algebra Lola Rodriguez Soalleiro

6:30 pm A didactical laboratory: image filtering A. Cavallone*, C. D’Apice*, M. Marsella & S. Salerno Computation of Normal Forms of Hamiltonian Systems in the Presence of Poisson Commuting integrals (Liouville-Integrability and Birkhoff Normal Forms) J. Mikram & F. Zinoun Busqueda automatica de lugares geometricos Eugenio Roanes Macias & Eugenio Roanes Lozano

June 26, 1999 (Saturday)

SESSION Computer Algebra meets Education II Bernhard Kutzler

A Computer Aided Geometric Design and Computer Algebra L. Gonzalez-Vega & J. R. Sendra

B Symbolic-Numeric Interface and Problem Solving Environments Richard Liska Stanly Steinberg & Robert van Engelen

C Computer Algebra for Dynamical Systems and Mechanics II Victor Edneral Ilias Kotsireas & Nikolay Vasiliev D Computations in Pure Mathematics (Algebra, Analysis, Geometry, ...) Maria-Emilia Alonso Francisco Castro & Laureano Gonzalez-Vega E

9:00 am A Tutorial Mathematical System for Elementary Schools Frank Postel* & Ralf Hillebrand Missing points and branches of real parametric curves Tomas Recio Finite Difference Numerical Modelling Supported by Computer Algebra Richard Liska Some properties of the symplectic Lie algebra Rodney Coleman Steenrod reduced powers and computability Rocio Gonzalez-Diaz & Pedro Real

9:30 am High-Powered Technology in a Low-Level Mathematics Course Jim Schultz Computational aspects of the N-body problem Ilias Kotsireas Analyzing the transference of the coalgebra structure on the homology of CDGAs M. J. Jimenez, P. Real & B. Silva

10:00 am A Scheme for Conservative Use of Computer Algebra Rein Prank Applying generalpurpose symbolic and numerical computation programs for Computer Graphics and CAGD. Some applications to industry Andres Iglesias & Akemi Galvez Ctadel: A Computer Algebra System for the Generation of Efficient Numerical Codes for PDEs Robert van Engelen The study by symbolic computation of the sign of the riemannian curvature in the configuration space of the 3body problem and applications B. Elmabsout & M. Barbosu Homology of semidirect product of groups V. Alvarez, J. A. Armario & P. Real

10:30 am Mathematica and didactical innovation: a quadric use case G. Albano*, A. Cavallone, C. D’Apice* & G. Gargiulo Proper Parametrization of Real Algebraic Surfaces Josef Schicho Invariant Variational Principles and Associated Numerical Schemes for Regularization of Ill-Posed Problems Ravi Venkatesan Oscillators in Resonance Antonio Elipe & Andr Deprit The use of Mathematica for the classification of some nilpotent Lie algebras L. M. Camacho, J. R. Gomez & R. M. Navarro

11:00 am  BREAK

11:30 am Computer Aided Instruction for Business Calculus in an International Liberal Arts College A. Kehagias & P. N. Vlachos* An algorithm for approximate conversion of rational B-spline curves/surfaces to integral B-spline curves/surfaces and its implementation Jesus Espinola, Laureano Gonzalez-Vega & Ioana Necula A Symbolic Numeric Environment for Analyzing Measurement Data in Multi-Model Settings Christoph Richard & Andreas Weber* Construction of Invariants of Simplectic Mapping Nikolay Vasiliev New ways of defining filiform Lie algebras J. C. Benjumea, F. J. Echarte, D. Fernandez, M. C. Marquez, J. Nuñez & F. Ramirez

12:00 pm Flexible Mathematical GUI Controls for Mathematical Electronical Documents Ralf Hillebrand* & Frank Postel Groebner Bases and Statistics Lorenzo Robbiano Group Invariant Finite-Difference Schemes for Advection Equation Ravi C. Venkatesan About an algorithm of T. Oaku Maribel Hartillo

12:30 pm Mathematica in Undergraduate Mathematics Courses—A Teaching or Learning Aid? May C. Abboud Usefulness of computer algebra methods in numerical simulations Michel Fournie Slopes in submodules of a free module Jose Maria Ucha

1:00 pm Searching roots for three degree polynomials with the help of Mathematica Marilo Lopez Gonzalez & Javier Rodrigo Hitos* Prototyping Symbolic-Numeric Algorithms using Naglink Brian J. Dupee* & James H. Davenport Computing Toric First Syzygies P. Pison-Casares & A. Vigneron-Tenorio

June 26, 1999 (Saturday)

SESSION Computer Algebra meets Education II Bernhard Kutzler

A Demos of Computer Algebra Systems Michael Wester & Winfried Neun

B Symbolic-Numeric Interface and Problem Solving Environments Richard Liska Stanly Steinberg & Robert van Engelen

C Applications of Computer Algebra to Robotics Peter Kovacs

D Computations in Pure Mathematics (Algebra, Analysis, Geometry, ...) Maria-Emilia Alonso Francisco Castro & Laureano Gonzalez-Vega E

3:30 pm Derivations and Visualizations for the Hydrogen Atom Joseph D. Myers*, Kelley B. Mohrmann & Terry T. Crow Some Noteworthy Issues Michael Wester A symbolicnumerical package for linear stability analysis of numerical methods for ODEs Massimo Cafaro & Beatrice Paternoster* Forward kinematics of parallel robots Jean-Pierre Merlet The ideal generation conjecture for s general rational curves F. Orecchia

4:00 pm CONVODE: A Reduce package for solving differential equations Alain Moussiaux Solving a RLC Circuit using Convolution with DERIVE for Windows Michel Beaudin Extraction of Low Order Boolean Rules from Trained Neural Networks using a Computer Algebra System Terence Etchells Using Computer algebra tools for off-line studies of parallel manipulators Jean-Charles Faugere, Luc Rolland & Fabrice Rouillier* Basic Algorithms for Specialization in Groebner Bases A. Montes

4:30 pm Exploring sequences with DERIVE Paolo Boieri The Computer Algebra System CoCoA Lorenzo Robbiano .

Self-Motions of Griffis-Duffy Type Parallel Manipulators Manfred Husty & Adolf Karger The Generation of Polya Polynomials using Macsyma, with Applications Bill Pletsch

5:00 pm  BREAK

SESSION Teaching of Efficient Mathematics A. Akritas & Gennadi Malaschonok Interaction on Physics, Number Theory & Computer Science I Hoang Ngoc Minh

5:30 pm A project of Computer Aided Learning: A course of Computer Algebra in the web Rafael J. Villanueva Mico & A. Hervas The REDLOG Package—Tools and Applications Volker Weispfenning Colored multiple zeta values relations table M. Bigotte Functional Ideal Decomposition—a Tool for Kinematics Peter Kovacs Duality in Effective Algebraic Geometry Bernard Mourrain

6:00 pm A study of particular methods for the approximate construction of some regular polygons by using Mathematica 3.0 Juan A. Aledo, Juan C. Cortes & Fernando L. Pelayo MuPAD—An Open Computer Algebra System and its Approach of Software Integration Ralf Hillebrand & Frank Postel Nested Sums and Iterated Integrals D. Bradley A Fast Computation of Hook Schur Functions F. Gaeta

6:30 pm Computer Algebra Applied to Mathematical Cartography Carlos Enriguez Turiño

evening  7:00pm - 8:00pm: Business Session in A

Program for the 5th International IMACS Conference on Applications of Computer Algebra

June 27, 1999 (Sunday)

SESSION Teaching of Efficient Mathematics A. Akritas & Gennadi Malaschonok

A Mathematics on the Internet Angel Diaz & Erich Kaltofen

B Applications of Quantifier Elimination Volker Weispfenning & Hoon Hong

C Interaction on Physics, Number Theory & Computer Science II Hoang Ngoc Minh

D

9:00 am A comparative study between two general methods for the approximate construction of regular polygons by using Mathematica 3.0 Juan A. Aledo, Juan C. Cortes & Fernando L. Pelayo Interfacing Proof Checkers With OpenMath Olga Caprotti A Robust Control System Design by a Special Quantifier Elimination Method using a Sturm-Habicht Sequence Hirokazu Anai & Shinji Hara Functional Equations on Polylogarithms with Axiom Hoang Ngoc Minh

9:30 am Possibilities for Alternative Uses of the ”Calculus and Mathematica” Approach Alkiviadis Akritas, Stelios Kapranidis & Athina Katalifou Plug and Play Mathematical Components Mike Dewar A Special Quantifier Elimination Algorithm for Pham Systems Laureano Gonzalez-Vega & Neila Gonzalez-Campos Algebraic Structures on the Set of Multiple Zeta Values Mike Hoffman

10:00 am Classical Mathematics with Mathematica Alkiviadis Akritas & Zamir Bavel To Be Announced Angel L. Diaz A new solution of the quartic problem: with application to integration David J. Jeffrey Computer aided Knot Theory using Mathematica and MathLink N. Imafuji

10:30 am Calculus and the Race Track Principle Alkiviadis Akritas & Zamir Bavel Mathematics on the Web— Past, Present and Future Dave Raggett Reach Set Computation for Linear Vector Fields using Quantifier Elimination Gerardo Lafferriere, George J. Pappas & Sergio Yovine Combinatorics on Algebraic Identification G. Jacob

11:00 am  BREAK

11:30 am Some elementary mathematics revisited and revised David J. Jeffrey Unicode Encoding of Mathematics Murray Sargent Quantifier Elimination for Trigonometric Polynomials by Cylindrical Trigonometric Decomposition Petru Pau & Josef Schicho Computational Construction of representation matrices for parallel version polynomial invariants M. Ochiai

12:00 pm Teaching Error-Correcting Codes, Discrete Mathematics and Modern Algebra with Computer Algebra Igor Gachkov & Kenneth Hulth Stylesheets for Mathematical Web Pages Stephen Watt The Study of Stability using Computer Algebra Stanly Steinberg Computational Decomposition of homeomorphisms into canonical Dehn’s twists Y. Sakata

12:30 pm Teaching of efficient mathematics Gennadi Malaschonok & Natalia Malaschonok Accessing Quantifier Elimination Packages Over the Internet Out of General Purpose Systems Andreas Weber On l-adic iterated integral Z. Wojtkowiak

1:00 pm Efficient methods of mathematical analysis Gennadi Malaschonok Concluding Discussion of the Issues Addressed in the Talks (Moderator with Audience Participation) Erich Kaltofen Semilinear motion planning in REDLOG Volker Weispfenning